What type of solution do you get for quadratic equations where D < 0? Give reasons for your answer. Also provide an example of such a quadratic equation and find the solution of the equation.

Answer by stanbon(75887)

What type of solution do you get for quadratic equations where D < 0? Give reasons for your answer. Also provide an example of such a quadratic equation and find the solution of the equation.
---------------
If the discriminant is negative the quadratic has two unequal complex
number solutions.
----------------
Example:
y = x^2 +x + 5
D = 1^2-4*1*5 = -19
x = [-1 +-sqrt(-19)]/2
x = [-1 + isqrt(19)]/2 or x = [-1 - isqrt(19)]/2
==========================
Cheers,
Stan H.